OpenStudy (anonymous):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2.

OpenStudy (anonymous):

@only_hanaaa @ganeshie8

OpenStudy (shamim):

f(g(x))=?

OpenStudy (shamim):

f(g(x))=f(-2x-3)=?

OpenStudy (shamim):

I just put the vakue of g(x)

OpenStudy (shamim):

f(-2x-3)=-(-2x-3)+4=?